# Improving accuracy of tree-tensor network approach by optimization of network structure

### v1

@@ III.2 Random XY-exchange Model Under Random Magnetic Field
Figure 7: Histogram $P(\mathcal{S})$ of entanglement entropy $\mathcal{S}$ on auxiliary bonds in TTNs obtained by the types where the initial TTN is (a) (1) MPN, (b) (2) PBT, and (c) (3) TTN obtained by tSDRG, for $\mathcal{N}_{\rm s}=200$ random samples. Circles connected by solid lines represent the data at $\chi=40$. Dashed (red), dot-dashed (green), and dotted (orange) lines are the data at $\chi=32,24$, and $16$, respectively.
@@
Those observations on the EEs are consistent with the findings obtained from the analyses of the variational energies. We thereby conclude that the stochastic structural optimizations [schemes (D) and (E)] work well for the random XY model and succeed in obtaining the best TTN structure yielding the lowest variational energy independently of the initial TTN.
@@ III.3 Richardson Model
We have applied the TTN structural optimization algorithm to the Richardson model, that is a spin-1/2 model containing the all-to-all XY-exchange interactions under the magnetic field. The Hamiltonian is given by
@@
so that the field $h_{i}$ linearly increases from $h_{1}=-\frac{1}{2}$ to $h_{N}=\frac{1}{2}$ with increasing the site index $i$. The Richardson model [Eq. (14)] is known to be exactly solvable and has been studied for many years in condensed-matter physics, particularly in connection with the BCS superconductivity, and nuclear physics.[33, 34, 35]
@@
As in the case of the random XY model in the previous section, we have performed several types of the structural optimization calculations to obtain the lowest-energy state in the subspace of zero total magnetization $\sum_{i}S^{z}_{i}=0$. The exchange constants and system sizes treated were $J_{\rm R}=1,4,16,64$ for $N=64$ and $J_{\rm R}=1$ for $N=128$. The maximum bond dimension treated was $\chi=80$. In the Richardson model, in which all spin pairs are coupled by the exchange interaction of the same strength, we can treat the system size $N$ and the bond dimension $\chi$ larger than those for the random XY model since we can carry out the calculation of exchange interaction terms with a much smaller (roughly speaking, $1/N$ times smaller) computational cost.
@@
We performed the calculation of schemes (A) - (E) starting from the initial TTN of (1) MPN and (2) PBT. For schemes (B)-(E), the structural optimization was carried out at $\chi=\chi_{\rm opt}=20$ or $40$, and then, the calculation to optimize only the isometries was done with increasing $\chi$ up to 80. See Appendix B for the details. For stochastic schemes (D) and (E), we performed the calculations of four runs, i.e., two runs (with different random seeds) for each of the annealing parameters $(\beta_{0},n_{\tau})=(0.1,2)$ and $(0.1,4)$. In addition, in order to avoid the deadlock of the calculation due to isometry whose all three legs have dimension $\chi=1$ discussed in Sec. II, we adopted an exceptional treatment in the preparation process of the initial TTN. The details of the treatment are presented in Appendix A. In our calculations, the treatment was necessitated only for the cases where $(J_{\rm R},N)=(64,64)$ and the initial TTN was PBT.
@@
Figure 8: Local spin polarization $\langle S^{z}_{i}\rangle$. The plotted are the results obtained by the type which yields the smallest relative error $\delta e$ for each model parameter $(J_{\rm R},N)$. Thick black line represents the magnetic field $h_{i}$.
@@
Before proceeding to the performance evaluation, let us examine the nature of the lowest energy state to be analyzed. In the Richardson model, the XY exchange interactions, which entangle all spins to form a large spin cluster, compete with the magnetic fields, which polarize the spins in the $S^{z}$ direction. Figure 8 presents the data of the spin polarization $\langle S^{z}_{i}\rangle$ in the lowest-energy states. For $J_{\rm R}=1$, when the exchange interaction is relatively weak, most of the spins except for those in the vicinity of $i=N/2$ are strongly polarized, $\langle S^{z}_{i}\rangle\sim\pm 1/2$, and almost classically ordered. Only the spins near $i=N/2$, where the field is weak, have the polarization slightly reduced by the exchange interactions; For example, $\langle S^{z}_{65}\rangle=-\langle S^{z}_{64}\rangle=0.4801$ for $J_{\rm R}=1$ and $N=128$. In this situation, the accuracy of the TTN calculation depends on how precisely the TTN state can represent the entanglement among those spins near $i=N/2$. As the exchange interaction $J_{\rm R}$ increases, the magnitude of the spin polarization decreases gradually, signaling that those spins start to entangle with each other to form a large spin cluster. A point to be noted here is that the $i$ dependence of the spin polarization remains flat except for the narrow region around $i=N/2$. Then, it is expected that the position in the TTN of those spins with nearly the same polarization does not affect the accuracy of the calculation very much. (Note that the exchange interactions take the same strength for all spin pairs and have no position dependence.) This point will be relevant in the later discussion on the TTN structure.
@@
In order to discuss the accuracy of the TTN optimization quantitatively, we explore the relative error in the variational energies defined by
$\delta e({\rm X},\chi)$ $=$ $\frac{E_{0}({\rm X},\chi)-E_{\rm exact}}{|E_{\rm exact}|},$ (16)
$\delta e^{\prime}({\rm X},\chi)$ $=$ $\frac{E^{\prime}_{0}({\rm X},\chi)-E_{\rm exact}}{|E_{\rm exact}|},$ (17)
@@
where $E_{\rm exact}$ is the exact lowest energy obtained from the exact solution of the Richardson model[33, 34, 35]; see Appendix C for the calculation of $E_{\rm exact}$. We will see that $\delta e$ and $\delta e^{\prime}$ exhibit qualitatively similar behavior, although $\delta e^{\prime}$ is by definition not less than $\delta e$. However, the difference between $\delta e$ and $\delta e^{\prime}$ is noticeable when the TTN is PBT [type (2A)]. This is because in PBT, the EE on the “top” bond is significantly larger than EEs on other bonds and $\delta e$ obtained at the top bond can avoid the effect of the truncation at the bond, which makes $\delta e$ lower than $\delta e^{\prime}$.
Figure 9: Relative error in the variational energy, $\delta e({\rm X},\chi)$, as functions of $\chi$, for the types where the initial TTN is (a) (1) MPN and (b) (2) PBT. The model parameters are $J_{\rm R}=1$ and $N=128$. For schemes (D) and (E), among the results of four runs, the ones of the run yielding the lowest variational energy at $\chi=80$ are plotted.
@@
Figure 10: Relative errors in the variational energies, $\delta e({\rm X},\chi)$ and $\delta e^{\prime}({\rm X},\chi)$, at (a) $\chi=60$ and (b) $\chi=80$. The model parameters are $J_{\rm R}=1$ and $N=128$. Circles and squares (triangles and diamonds) represent $\delta e({\rm X},\chi)$ [$\delta e^{\prime}({\rm X},\chi)$] obtained by the types where the initial TTN is (1) MPN and (2) PBT, respectively. Solid symbols represent the results of (A) or (B)-(E) with $\chi_{\rm opt}=20$, while open symbols are those of (B)-(E) with $\chi_{\rm opt}=40$. For schemes (D) and (E), the results of four runs are plotted. Downward triangles represent $\delta e({\rm X},\chi)$ obtained by the calculations of schemes (B) and (C) with $\chi_{\rm opt}=40$ starting from PBTs with four different random site orderings (rso); see the main text. Some of the results for schemes (D) and (E) and the initial PBTs with random-site orderings overlap.
@@
Let us compare $\delta e$ ($\delta e^{\prime}$) of various types of the calculation. Figure 9 shows the $\chi$ dependence of $\delta e({\rm X},\chi)$ for $(J_{\rm R},N)=(1,128)$. Also shown in Fig. 10 are the data of $\delta e({\rm X},\chi)$ and $\delta e^{\prime}({\rm X},\chi)$ for $\chi=60$ and $80$. We find that for the calculations with a fixed TTN structure [scheme (A)], PBT gives a rather accurate result. The non-stochastic optimization schemes [schemes (B) and (C)] yield results with better accuracy than those of scheme (A), except for (1B), whose accuracy is more or less the same as that of (1A). In particular, scheme (C) succeeded in realizing a sizable reduction of the relative error at $\chi=60$, compared to scheme (A) [see Fig. 10(a)]. The relative error $\delta e$ for (1C) with $\chi_{\rm opt}=20$ at $\chi=60$ and $80$ converge up to $\delta e(\chi=60)=9.3\times 10^{-12}$ and $\delta e(\chi=80)=8.8\times 10^{-12}$, indicating that the calculation gives a nearly exact result that reaches the accuracy limit of our TTN variational calculation (with double precision variables and the rule to discard the bases in isometry with singular values smaller than $10^{-6}$). These observations suggest that the non-stochastic structural optimization schemes are able to realize the accuracy improvement for the Richardson model, although the improvement may occasionally be small, like (1B) of the present case. On the other hand, the accuracy of the calculations of stochastic structural optimization [schemes (D) and (E)] is poor, even worse than that without structural optimization [scheme (A)], indicating that they do not work well for the Richardson model.
@@
Figure 11: Relative errors in the variational energies, $\delta e({\rm X},\chi)$ and $\delta e^{\prime}({\rm X},\chi)$, at $\chi=80$ for $N=64$ and (a) $J_{\rm R}=4$, (b) $J_{\rm R}=16$, and (c) $J_{\rm R}=64$. Circles and squares (triangles and diamonds) represent $\delta e({\rm X},\chi)$ [$\delta e^{\prime}({\rm X},\chi)$] obtained by the types where the initial TTN is (1) MPN and (2) PBT, respectively. Solid symbols represent the data of (A) or (B)-(E) with $\chi_{\rm opt}=20$, while open symbols are those of (B)-(E) with $\chi_{\rm opt}=40$. For schemes (D) and (E), the results of four runs are plotted. Downward triangles represent $\delta e({\rm X},\chi)$ obtained by the calculations of schemes (B) and (C) with $\chi_{\rm opt}=40$ starting from PBTs with four different random site orderings (rso); see the main text. Some of the results for schemes (D) and (E) and the initial PBTs with random-site orderings overlap.
@@
In Fig. 11, we show the results of $\delta e$ and $\delta e^{\prime}$ for $J_{\rm R}=4,16,64$ and $N=64$. These data for large $J_{\rm R}$ also indicate basically the same results as those for $J_{\rm R}=1$. That is, when the initial TTN is MPN, the non-stochastic optimization [schemes (B) and (C)] may achieve accuracy improvement, although the calculation is occasionally stuck in a state of poor accuracy. PBT yields a result with good accuracy as it is, but even better accuracy may be achieved by the calculation of scheme (C). In contrast, the calculations of stochastic schemes (D) and (E) are not accurate and do not work well for the Richardson model.
@@
Figure 12: (a) Average value $\mathcal{S}_{\rm av}$ and (b) maximum value $\mathcal{S}_{\rm max}$ of EEs on auxiliary bonds for $J_{\rm R}=1$ and $N=128$. Circles and squares represent the data at $\chi=80$ of the types where the initial TTN is (1) MPN and (2) PBT, respectively. Solid symbols represent the data of (A) or (B)-(E) with $\chi_{\rm opt}=20$, while open symbols are those of (B)-(E) with $\chi_{\rm opt}=40$. For schemes (D) and (E), the results of four runs are plotted. The $\chi$ convergence is confirmed by comparing the results at $\chi=80$ and $60$. Some of the results for schemes (D) and (E) overlap.
@@
Figure 12 shows the average and maximum values of the EE on auxiliary bonds, $\mathcal{S}_{\rm av}$ and $\mathcal{S}_{\rm max}$, for $J_{\rm R}=1$ and $N=128$. The average value $\mathcal{S}_{\rm av}$ is remarkably large for (1A), where the TTN structure is fixed to be MPN, while $\mathcal{S}_{\rm av}$ for the other types are small. The maximum values $\mathcal{S}_{\rm max}$ for (1A) and (2A), where the TTN structures are MPN and PBT, are the same and large, which are the values on the bond corresponding to the bipartition of the system into two parts of $21\leq i\leq N/$ 2 and $N/2+1\leq i\leq N$. $\mathcal{S}_{\rm max}$ for other types are smaller than those of (1A) and (2A); in particular, scheme (B) yields the smallest $\mathcal{S}_{\rm max}$, indicating that scheme (B), which directly uses the EE as a measure in selecting the local TTN structure, is the most suitable for minimizing the maximum value of EE. The qualitatively same features of $\mathcal{S}_{\rm av}$ and $\mathcal{S}_{\rm max}$ are also observed for other model parameters $J_{\rm R}$ and $N$ (data not shown), while the values of $\mathcal{S}_{\rm av}$ and $\mathcal{S}_{\rm max}$ increase with $J_{\rm R}$. Here, we recall that the relative errors in the variational energy were smallest for scheme (C) in most cases. These results thus suggest that there was no strong correlation between the accuracy of the variational energy and the average and maximum EEs in the Richardson model. The observations indicate the importance of choosing the adequate scheme of the network optimization depending on the model treated and the quantity to be minimized.
@@
Figure 13: Histogram $P(\mathcal{S})$ of entanglement entropy $\mathcal{S}$ on auxiliary bonds in TTNs obtained by the types where the initial TTN is (a) (1) MPN and (b) (2) PBT for $J_{\rm R}=1$ and $N=128$. Circles connected by solid lines represent the data at $\chi=80$. The $\chi$ convergence is confirmed by comparing the results at $\chi=80$ and $60$. The size of bins is set to be $\Delta\mathcal{S}=0.02$. For schemes (D) and (E), among the results of four runs, the one obtained by the run yielding the lowest variational energy at $\chi=80$ is plotted.
@@
The histograms of EEs on auxilialy bonds for $(J_{\rm R},N)=(1,128)$ are shown in Fig. 13. The bin size is set to be $\Delta\mathcal{S}=0.02$. We find that only the histogram of type (1A), which uses the TTN structure fixed to MPN, exhibits a slowly decreasing behavior that is different from the other types. The histograms except the one of (1A) show a similar structure in which the frequency takes the largest value at a small EE and rapidly decreases as the EE increases. The histograms for the other values of $J_{\rm R}$ and $N$ (not shown here) exhibit qualitatively the same behavior, while the scale of the horizontal axis (the values of EE) increases with $J_{\rm R}$. It is thus suggested that in the Richardson model, the histograms of EE also have no strong correlation with the accuracy in the variational enegy.
@@
Figure 14: TTN structure obtained by type (2C) with $\chi_{\rm opt}=40$, which yielded the lowest variational energy at $\chi=80$, for $J_{\rm R}=1$ and $N=128$. Open circles represent isometries and the solid diamond represents the singular-value matrix. Bare spins are represented by circles with colors varying according to the site index $i$ as shown in the color bar. Bare spins with the site index $57\leq i\leq 72$ are indicated by numbers.
@@
Figure 15: TTN structure obtained by a calculation of type (2D) with $\chi_{\rm opt}=40$, which yielded a high vatiational energy, for $J_{\rm R}=1$ and $N=128$. Open circles represent isometries and the solid diamond represents the singular-value matrix. Bare spins are represented by circles with colors varying according to the site index $i$ as shown in the color bar. Bare spins with the site index $57\leq i\leq 72$ are indicated by numbers.
@@
Finally, we examine the TTN structure obtained. As discussed above, the optimal TTN for the Richardson model with Eq. (15) is expected to have a structure suitable for representing the entanglement among quantum-fluctuating spins around $i=N/2$. Figure 14 presents the TTN structure obtained by the calculation of type (2C) with $\chi_{\rm opt}=40$ for $J_{\rm R}=1$ and $N=128$, which yielded the lowest variational energy. In this TTN, the spins around $i=N/2$ are located close to each other, indicating that a structure nearly the optimal one is realized by the calculation. (To be precise, as the spins at the sites $69\leq i\leq 72$ are not far away but slightly apart from the clusters of the spins at $57\leq i\leq 68$ in the TTN obtained, the TTN should not be the “truly optimal”, although it is close enough to the optimal to yield the lowest variational energy.) Note that in the initial PBT of the present calculation, the spins around $i=N/2$ are separated into the two groups of $i\leq N/2$ and $i\geq N/2+1$, which are located far apart; therefore, the initial PBT is not suitable for representing the entanglement between the groups, although entanglement among the spins within each group can be taken into account well. Our algorithm achieves a structural change from the initial PBT to the nearly optimal TTN. We have also confirmed that the TTNs with the smallest variational energy obtained for other model parameters, $J_{\rm R}=1,4,16,64$ and $N=64$ have the similar structure in which the spins near $i=N/2$ are clustered close together, as in Fig. 14.

### v2

@@ III.2 Random XY-exchange Model Under Random Magnetic Fields
Figure 7: Histogram $P(\mathcal{S})$ of entanglement entropy $\mathcal{S}$ on auxiliary bonds in TTNs obtained by the calculations where the initial TTN is (a) MPN, (b) PBT, and (c) tSDRG-TTN, for $\mathcal{N}_{\rm s}=200$ random samples. “NoOpt” represents the calculation without structural optimization, whereas “bondEE”, “rDMwgt”, and “Stoch” represent respectively the optimization using the bond EE, that using the truncated rDM weight, and the stochastic optimization. Circles connected by solid lines represent the data at $\chi=40$. Dashed (red), dot-dashed (green), and dotted (orange) lines are the data at $\chi=32,24$, and $16$, respectively. The size of bins is set to $\Delta\mathcal{S}=0.1$.
We also analyze the histogram of EEs obtained by each type of calculation. We calculate the histogram $P(\mathcal{S})$ using the EEs on the auxiliary bonds in all random samples [the total number of EEs is $\mathcal{N}_{\rm s}(N-3)$], where the bin size is set to $\Delta\mathcal{S}=0.1$. Figure 7 presents the obtained histograms. The $\chi$ dependence is sizable only in the histogram of the calculation without structural optimization for MPN and irrelevant for the other results. The histograms for the scheme without structural optimization have a different shape depending on the initial TTNs, which should reflect the characteristics of each TTN. When applied to PBT, the non-stochastic optimizations yield histograms with a similar shape to that of the calculation without structural optimization, suggesting that the TTN structures obtained there remain under the influence of the initial PBT. Also, when the initial TTN is MPN or tSDRG-TTN, the histograms obtained by the non-stochastic optimizations still exhibit peculiar characteristics such as bumpy behavior in the region of large EE ($1.2\lesssim\mathcal{S}\lesssim 3.0$), although their overall structures tend to approach the histograms obtained by the stochastic optimization. In contrast, the histograms obtained by the stochastic optimization converge to nearly the same shape regardless of the initial TTN; $P(\mathcal{S})$ exhibits a sharp peak around $\mathcal{S}\sim 1.0-1.1$, and from the peak, it decreases with a gentle tail as $\mathcal{S}$ increases. These histograms should be the ones of the best TTNs achieved in the present results. Finally, no essential difference is found between the results for $\chi_{\rm opt}=16$ and $32$.
@@
Those observations on the EEs are consistent with the findings obtained from the analyses of the variational energies. We thereby conclude that the stochastic structural optimization works well for the random XY model and succeeds in obtaining the best TTN structure yielding the lowest variational energy independently of the initial TTN.
@@ III.3 Richardson Model
We have applied the TTN structural optimization algorithm to the Richardson model, that is a spin-1/2 model containing all-to-all XY-exchange interactions under magnetic fields. The Hamiltonian is given by
@@
so that the field $h_{i}$ linearly increases from $h_{1}=-\frac{1}{2}$ to $h_{N}=\frac{1}{2}$ with the site index $i$. The Richardson model [Eq. (14)] is known to be exactly solvable and has been studied for many years in condensed-matter physics, particularly in connection with the BCS superconductivity, and nuclear physics.[33, 34, 35]
@@
As in the case of the random XY model in the previous section, we have performed several types of the structural optimization calculations to obtain the lowest-energy state in the subspace of zero total magnetization $\sum_{i}S^{z}_{i}=0$. The exchange constant and system size were $J_{\rm R}=1,4,16,64$ for $N=64$ and $J_{\rm R}=1$ for $N=128$. The maximum bond dimension was $\chi=80$. Here, we note that the system size $N$ and the bond dimension $\chi$ are much larger than those for the random XY model. This is because the computational cost of the exchange interaction terms for the Richardson model is much smaller (roughly speaking, $1/N$ times smaller) than that for the random XY model, reflecting the property of the Richardson model that all spin pairs are coupled by the exchange interaction of the same strength.
@@
We have performed the calculations starting from the initial TTN of MPN and PBT. In the calculations without structural optimization, the bond dimension was increased from $\chi=20$ to $80$. In the calculations with structural optimization, the TTN structure was optimized at $\chi=\chi_{\rm opt}=20$ or $40$, and then, the calculation to optimize only the isometries was done with $\chi$ up to 80. See Appendix B for the details. For the stochastic optimization, we have performed the calculations of four runs, i.e., two runs (with different random seeds) for each set of the annealing parameters $(\beta_{0},n_{\tau})=(0.1,2)$ and $(0.1,4)$. In addition, in order to avoid the deadlock of the calculation due to an isometry whose all three legs have dimension $\chi=1$ discussed in Sec. II, we adopted an exceptional treatment in the preparation process of the initial TTN. The details of the treatment are presented in Appendix A. In our calculations, the treatment was necessitated only for the cases where $(J_{\rm R},N)=(64,64)$ and the initial TTN was PBT.
@@
Figure 8: Local spin polarization $\langle S^{z}_{i}\rangle$. The plots show the results obtained by the calculation that yields the smallest relative error $\delta e$ for each model parameter $(J_{\rm R},N)$. The thick black line represents the magnetic field $h_{i}$.
@@
Before proceeding to the performance evaluation, let us examine the nature of the lowest energy state to be analyzed. In the Richardson model, the XY exchange interactions, which entangle all spins to form a large spin cluster, compete with the magnetic fields, which polarize the spins in the $S^{z}$ direction. Figure 8 presents the data of the spin polarization $\langle S^{z}_{i}\rangle$ in the lowest-energy states. For $J_{\rm R}=1$, where the exchange interaction is relatively weak, most of the spins except for those in the vicinity of $i=N/2$ are strongly polarized, $\langle S^{z}_{i}\rangle\sim\pm 1/2$, and almost classically ordered. Only the spins near $i=N/2$, where the field is weak, have the polarization slightly reduced by the exchange interactions; For example, $\langle S^{z}_{65}\rangle=-\langle S^{z}_{64}\rangle=0.4801$ for $J_{\rm R}=1$ and $N=128$. In this situation, the accuracy of the TTN calculation depends on how precisely the TTN state can represent the entanglement among those spins near $i=N/2$. As the exchange interaction $J_{\rm R}$ increases, the magnitude of the spin polarization decreases gradually, signaling that the spins start to entangle with each other to form a large spin cluster. A point to be noted here is that the $i$ dependence of the spin polarization remains flat except for the narrow region around $i=N/2$. Thus, it is expected that the location of such spins with nearly the same polarization in a TTN is irrelevant to the accuracy of the calculation. (Note that the exchange interactions take the same strength for all spin pairs and have no position dependence.) This point will be essential in the later discussion on the TTN structure.
@@
In order to discuss the accuracy of the TTN optimization quantitatively, we evaluate the relative error in the variational energies defined by
$\delta e(\chi)$ $=$ $\frac{E_{0}(\chi)-E_{\rm exact}}{|E_{\rm exact}|},$ (16)
$\delta e^{\prime}(\chi)$ $=$ $\frac{E^{\prime}_{0}(\chi)-E_{\rm exact}}{|E_{\rm exact}|},$ (17)
@@
where $E_{\rm exact}$ is the exact lowest energy obtained from the exact solution for the Richardson model[33, 34, 35]; See Appendix D for the calculation of $E_{\rm exact}$. We will see that $\delta e$ and $\delta e^{\prime}$ exhibit qualitatively similar behavior, although $\delta e^{\prime}$ is by definition not less than $\delta e$. However, the difference between $\delta e$ and $\delta e^{\prime}$ is noticeable for PBT without structural optimization. This is because in PBT, the EE on the “top” bond is significantly larger than EEs on other bonds, and $\delta e$ calculated at the top bond can be further reduced by avoiding the effect of the truncation of that large EE.
Figure 9: Relative error in the variational energy, $\delta e(\chi)$, for the calculations where the initial TTN is (a) MPN and (b) PBT. The model parameters are $J_{\rm R}=1$ and $N=128$. “NoOpt” represents the calculation without structural optimization, whereas “bondEE”, “rDMwgt”, and “Stoch” represent respectively the optimization using the bond EE, that using the truncated rDM weight, and the stochastic optimization. For the stochastic optimization, the result of the run yielding the lowest variational energy at $\chi=80$ among the four runs is plotted.
@@
Figure 10: Relative errors in the variational energies, $\delta e(\chi)$ and $\delta e^{\prime}(\chi)$, at (a) $\chi=60$ and (b) $\chi=80$. The model parameters are $J_{\rm R}=1$ and $N=128$. Circles and squares (triangles and diamonds) represent $\delta e(\chi)$ $\delta e^{\prime}(\chi)$] for the calculations where the initial TTN is MPN and PBT, respectively. “NoOpt” represents the calculation without structural optimization, whereas “bondEE”, “rDMwgt”, and “Stoch” represent respectively the optimization using the bond EE, that using the truncated rDM weight, and the stochastic optimization. For the calculations with structural optimization, the results for $\chi_{\rm opt}=20$ ($40$) are shown by solid (open) symbols. For the stochastic optimization, the results of four runs are plotted. Downward triangles represent $\delta e(\chi)$ obtained by the calculations of non-stochastic structural optimizations with $\chi_{\rm opt}=40$ starting from PBTs with four different random site orderings (rso); See the main text. For the stochastic optimization and the non-stochastic optimizations with the initial PBTs with random-site orderings, some of the results overlap.
@@
Let us compare $\delta e$ ($\delta e^{\prime}$) of various types of calculation. Figure 9 shows the $\chi$ dependence of $\delta e(\chi)$ for $(J_{\rm R},N)=(1,128)$. We also show in Fig. 10 the data of $\delta e(\chi)$ and $\delta e^{\prime}(\chi)$ for $\chi=60$ and $80$. We find that for the calculations without structural optimization, PBT gives a rather accurate result. As for the non-stochastic optimization schemes, the optimization using the truncated rDM weight realizes a sizable reduction in the relative error; $\delta e$ obtained by the optimization for the initial MPN with $\chi_{\rm opt}=20$ converge up to $\delta e(\chi=60)=9.3\times 10^{-12}$ at $\chi=60$ and $\delta e(\chi=80)=8.8\times 10^{-12}$ at $\chi=80$, indicating that the calculation even at $\chi=60$ gives a nearly exact result within the accuracy limit of our calculation.[36] On the other hand, the optimization using the bond EE reduces the relative error compared to that without structural optimization when applied to the initial PBT, while the reduction is almost negligible for the initial MPN. These observations suggest that the non-stochastic structural optimization schemes are able to realize the accuracy improvement for the Richardson model, although the improvement may occasionally be small. In contrast, the accuracy of the stochastic structural optimization is poor, even worse than that without structural optimization, indicating that the stochastic optimization does not work well for the Richardson model.
@@
Figure 11: Relative errors in the variational energies, $\delta e(\chi)$ and $\delta e^{\prime}(\chi)$, at $\chi=80$ for $N=64$ and (a) $J_{\rm R}=4$, (b) $J_{\rm R}=16$, and (c) $J_{\rm R}=64$. Circles and squares (triangles and diamonds) represent $\delta e(\chi)$ $\delta e^{\prime}(\chi)$] for the calculations where the initial TTN is MPN and PBT, respectively. “NoOpt” represents the calculation without structural optimization, whereas “bondEE”, “rDMwgt”, and “Stoch” represent respectively the optimization using the bond EE, that using the truncated rDM weight, and the stochastic optimization. For the calculations with structural optimization, the results for $\chi_{\rm opt}=20$ ($40$) are shown by solid (open) symbols. For the stochastic optimization, the results of four runs are plotted. Downward triangles represent $\delta e(\chi)$ obtained by the calculations of non-stochastic structural optimizations with $\chi_{\rm opt}=40$ starting from PBTs with four different random site orderings (rso); See the main text. For the stochastic optimization and the non-stochastic optimizations with the initial PBTs with random-site orderings, some of the results overlap.
@@
In Fig. 11, we show the results of $\delta e$ and $\delta e^{\prime}$ for $J_{\rm R}=4,16,64$ and $N=64$. These data for large $J_{\rm R}$ also exhibit basically the same results as those for $J_{\rm R}=1$. That is, when the initial TTN is MPN, the non-stochastic optimizations can achieve accuracy improvement, although they are occasionally stuck in a state of poor accuracy. PBT yields a result with good accuracy as it is, but even better accuracy may be achieved by the optimization using the truncated rDM weight. In contrast, the stochastic optimization is not accurate and does not work well for the Richardson model.
@@
Figure 12: (a) Average value $\mathcal{S}_{\rm av}$ and (b) maximum value $\mathcal{S}_{\rm max}$ of EEs on auxiliary bonds for $J_{\rm R}=1$ and $N=128$. Circles and squares represent the results at $\chi=80$ of the calculations where the initial TTN is MPN and PBT, respectively. “NoOpt” represents the calculation without structural optimization, whereas “bondEE”, “rDMwgt”, and “Stoch” represent respectively the optimization using the bond EE, that using the truncated rDM weight, and the stochastic optimization. For the calculations with structural optimization, the results for $\chi_{\rm opt}=20$ ($40$) are shown by solid (open) symbols. For the stochastic optimization, the results of four runs are plotted. The $\chi$ convergence has been confirmed by comparing the results at $\chi=80$ and $60$. For the stochastic optimization, some of the results overlap.
@@
Figure 12 shows the average and maximum values of the EEs on auxiliary bonds, $\mathcal{S}_{\rm av}$ and $\mathcal{S}_{\rm max}$, for $J_{\rm R}=1$ and $N=128$. The average value $\mathcal{S}_{\rm av}$ is remarkably large for the calculation without structural optimization for the initial MPN, while $\mathcal{S}_{\rm av}$ is small for the other cases. The maximum values $\mathcal{S}_{\rm max}$ for the calculations without structural optimization, where the TTN structures are MPN or PBT, are equal and large, which are the EE between the subsystems of $1\leq i\leq N/2$ and $N/2+1\leq i\leq N$. Compared with them, $\mathcal{S}_{\rm max}$ for the other cases are small; In particular, the optimization using the bond EE yields the smallest $\mathcal{S}_{\rm max}$, indicating that the scheme, which directly uses the EE as a measure in selecting the local TTN structure, is the most suitable for minimizing $\mathcal{S}_{\rm max}$. The qualitatively same features of $\mathcal{S}_{\rm av}$ and $\mathcal{S}_{\rm max}$ are also observed for other model parameters $J_{\rm R}$ and $N$ (data not shown), while the values of $\mathcal{S}_{\rm av}$ and $\mathcal{S}_{\rm max}$ increase as $J_{\rm R}$ increases. Here, we recall that the relative errors in the variational energy were smallest for the optimization using the truncated rDM weight in most cases. These results thus suggest that there was no strong correlation between the accuracy of the variational energy and the average and maximum EEs in the Richardson model. The observations indicate the importance of choosing an adequate scheme of network optimization depending on the model treated and the quantity to be minimized.
@@
Figure 13: Histogram $P(\mathcal{S})$ of entanglement entropy $\mathcal{S}$ on auxiliary bonds in TTNs obtained by the calculations where the initial TTN is (a) MPN and (b) PBT for $J_{\rm R}=1$ and $N=128$. “NoOpt” represents the calculation without structural optimization, whereas “bondEE”, “rDMwgt”, and “Stoch” represent respectively the optimization using the bond EE, that using the truncated rDM weight, and the stochastic optimization. Circles connected by solid lines represent the data at $\chi=80$. The $\chi$ convergence has been confirmed by comparing the results at $\chi=80$ and $60$. The size of bins is set to $\Delta\mathcal{S}=0.02$. For the stochastic optimization, the result of the run yielding the lowest variational energy at $\chi=80$ among the four runs is plotted.
@@
The histograms of EEs on auxiliary bonds for $(J_{\rm R},N)=(1,128)$ are shown in Fig. 13. The bin size is set to be $\Delta\mathcal{S}=0.02$. We find that the histogram obtained by the calculation without structural optimization for the initial MPN exhibits a unique stairway-like distribution. The other histograms show a similar structure in which the frequency takes the largest value at a small EE and rapidly decreases as the EE increases. The histograms for the other values of $J_{\rm R}$ and $N$ (not shown here) exhibit qualitatively the same behavior, while the scale of the horizontal axis (the values of EE) increases with $J_{\rm R}$. It is thus suggested that in the Richardson model, the histograms of EE also have no strong correlation with the accuracy in the variational energy.
@@
Figure 14: TTN structure obtained by the optimization using the truncated rDM weight for the initial PBT with $\chi_{\rm opt}=40$, which yielded the lowest variational energy at $\chi=80$, for $J_{\rm R}=1$ and $N=128$. Open circles represent isometries and the solid diamond represents the singular-value matrix. Bare spins are represented by circles with colors varying according to the site index $i$ as shown in the color bar. Bare spins with the site index $57\leq i\leq 72$ are indicated by numbers.
@@
Figure 15: TTN structure obtained by a calculation of the stochastic optimization for the initial PBT with $\chi_{\rm opt}=40$, which yielded a high variational energy, for $J_{\rm R}=1$ and $N=128$. Open circles represent isometries and the solid diamond represents the singular-value matrix. Bare spins are represented by circles with colors varying according to the site index $i$ as shown in the color bar. Bare spins with the site index $57\leq i\leq 72$ are indicated by numbers.
@@
Finally, we examine the TTN structure obtained. As discussed above, the optimal TTN for the Richardson model with Eq. (15) is expected to have a structure suitable for representing the entanglement among quantum-fluctuating spins around $i=N/2$. Figure 14 presents the TTN structure obtained by the structural optimization using the truncated rDM weight for the initial PBT with $\chi_{\rm opt}=40$ for $J_{\rm R}=1$ and $N=128$, which yielded the lowest variational energy. In this TTN, the spins around $i=N/2$ are located close to each other, indicating that a structure nearly the optimal one is realized. (To be precise, as the spins at the sites $69\leq i\leq 72$ are not far away but slightly apart from the clusters of the spins at $57\leq i\leq 68$ in the TTN obtained, the TTN should not be the “truly optimal”, although it is close enough to the optimal to yield the lowest variational energy.) Note that in the initial PBT of the present calculation, the spins around $i=N/2$ are separated into the two groups of $i\leq N/2$ and $i\geq N/2+1$, which are located far apart; Therefore, the initial PBT is not suitable for representing the significant entanglement across the two groups. Our algorithm achieves a structural change from the initial PBT to the nearly optimal TTN. We have also confirmed that the TTNs with the smallest variational energy obtained for other model parameters, $J_{\rm R}=1,4,16,64$ and $N=64$ have a similar structure in which the spins near $i=N/2$ are clustered close together, as in Fig. 14.